OpenStudy (anonymous):

guys help , please? RoTo = RbTb + 16 Ro+Rb = 1 To= 1 TB= -6 Find Ro and Rb. i have no idea how to solve these!

OpenStudy (anonymous):

the answers are A) Ro = 3 Rb= -2 B) Ro=-3 Rb=2 C) Ro=2 Rb=-3 D) Ro=-2 Rb=3 E) none of these

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

@MLeighW

OpenStudy (anonymous):

First fill in the values you know.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so what do you have?

OpenStudy (anonymous):

Ro = -6Rb + 16

OpenStudy (anonymous):

Okay I am going to try this another way haha I am making this too complicated. Work backwards. Fill in you Ro and Rb with the values the answers and see what makes the statement true.

OpenStudy (anonymous):

alright any guesses u have?

OpenStudy (anonymous):

it is D right jst calculated

OpenStudy (anonymous):

So -3(1)=-3 obviously. 2(-6)=-12, -12+16=4 so it isn't a.

OpenStudy (anonymous):

2(1)=2 duhh lol -3(-6)=18 and 18+16=34. Not it once again...

OpenStudy (anonymous):

D ? does that work

OpenStudy (anonymous):

D there u go @owlgirl98 there's ur ans

OpenStudy (anonymous):

you tell me what you think. Try working it out. :)

OpenStudy (anonymous):

I did i think it's correct